# Set Features for Anomaly Detection

### v3

@@ 4.1 Images as Sets
Images can be seen as consisting of a set of elements of different levels of granularity. This ranges from pixels to small patches, to low-level elements such as lines or corners, up to high-level elements such as objects. For anomaly detection, we typically do not know in advance the correct level of granularity for separating between normal and anomalous samples [26]. This depends on the anomalies, which are unknown during training. Instead, we first use multiple levels of granularity, describing image patches of different sizes, and combine their scores.
In practice, we use representations from intermediate blocks of a pre-trained ResNet [25]. As a ResNet network simultaneously embeds many local patches of each image, we pass the image samples through the network encoder and extract our representations from the intermediate activations at the end of different ResNet blocks (see Fig.3). We define each spatial location in the activation map as an element. Note that as different blocks have different resolutions, they yield different numbers of elements per layer. We run our set methods with the elements at the end of each residual block used and combine the results in an ensemble as detailed in the supplementary material (App.13.1).
@@ 4.2 Time Series as Sets
Time series data can be viewed as a set of temporal windows. Similarly to images, it is generally not known in advance which temporal scale is relevant for detecting anomalies; i.e., what is the duration of windows which includes the semantic phenomenon. Inspired by Rocket [18], we define the basic elements of a time series as a collection of temporal window pyramids. Each pyramid contains $L$ windows. All the windows in a pyramid are centered at the same time step, each containing $\tau$ samples (Fig.3). The first level window includes $\tau$ elements with stride $1$, the second level window includes $\tau$ elements with stride $2$, etc. Such window pyramid is computed for each time step in the series. The entire series is represented as a set of pyramids of its elements. Implementation details for both modalities are described in Sec.13.2.
@@ 5.1 Logical Anomaly Detection Results
Logical Anomalies Dataset. We use the recently published MVTec-LOCO dataset [5] to evaluate our method’s ability to detect anomalies caused by unusual configurations of normal elements. This dataset features five different classes: breakfast box, juice bottle, pushpins, screw bag and splicing connector (see Fig.1). Each class includes: (i) a training set of normal samples ($\sim 350$ samples). (ii) a validation set, containing a smaller set of normal samples ($\sim 60$ samples). (iii) a test set, containing normal samples, structural anomalies, and logical anomalies ($\sim 100$ each).
The anomalies in each class are divided into structural anomalies and logical anomalies. Structural anomalies feature local defects, somewhat similar to previous datasets such as [3]. Conversely, logical anomalies may violate ‘logical’ conditions expected from the normal data. As one example, an anomaly may include a different number of objects than the numbers expected from a normal sample (while all the featured object types exist in the normal class Fig.1). Other types of logical anomalies in the dataset may include cases where distant parts of an image must correlate with one another. For instance, within the normal data, the color of one object may correlate with the length of another object. These correlations may break in an anomalous sample.
Baselines. We compare to baseline methods used by the paper which presented the MVTec-LOCO dataset [5]: Variational Model (VM) [57], MNAD, f-AnoGAN [55], AE / VAE. Student Teacher (ST), SPADE, PatchCore (PCore) [48]. We also compare to GCAD [5] - a reconstruction-based method, based on both local and global deep ResNet features, which was explicitly designed for logical anomaly detection. EfficientAD - [1], a reconstruction-based method, with a loss aimed at preventing an autoencoder from reconstructing well unseen images. We also report the results by PUAD [58], an ensemble method combining [1] and [47]. Finally, we report SINBAD+EfficientAD, a simple average of our and EfficientAD’s per-sample results. As the last set of baselines does not always report per-class accuracies, we report them in a different table.
@@
Results. We report per-class results on image-level detection of logical anomalies and structural anomalies in Tab.1. Interestingly, we find complementary strengths between our approach and GCAD, a reconstruction-based approach by [5]. Although GCAD performed better on specific classes (e.g., pushpins), our approach provides better results on average. Notably, our approach provides non-trivial anomaly detection capabilities on the screw bag class, while baseline approaches are close to the random baseline. EfficientAD [1], focuses on structural anomalies and achieves impressive results on them, but underperforms on logical anomalies (see Tab.2).
As our method is comparatively strong on specific classes (e.g., Screw bag Logical, $81.1\%$ compared to $56.0\%$ of GCAD and $55.5\%$ of EfficientAD), it makes a strong contributor to ensemble methods. For example, a simple combination of our method with EfficientAD [1] (SINBAD+EfficientAD) outperforms all other methods and ensembles (See Tab. 2).
Our approach also provides an improvement over detection-by-segmentation in the detection of structural anomalies in some classes. This is somewhat surprising, as one may assume that detection-by-segmentation approaches would perform well in these cases. One possible reason for that is the high variability of the normal data in some of the classes (e.g., breakfast box, screw bag, Fig.1). This high variability may induce false positive detections for baseline approaches. Taken together, while different methods provide complementary strengths, on average, our method provides state-of-the-art results in logical anomaly detection. See also the discussion at Sec.6
@@
See Tab.7 for the full table. $\sigma$ presented in Tab. 8
@@
5.2 Time series anomalies detection results
Time series dataset. We compared on the five datasets used in NeurTraL-AD [42]: RacketSports (RS). Accelerometer and gyroscope recording of players playing different racket sports. Each sport is designated as a class. Epilepsy (EPSY). Accelerometer recording of healthy actors simulating four activity classes, e.g. an epileptic shock. Naval air training and operating procedures standardization (NAT). Positions of sensors mounted on body parts of a person performing activities. There are six different activity classes in the dataset. Character trajectories (CT). Velocity trajectories of a pen on a WACOM tablet. There are $20$ characters in this dataset. Spoken Arabic Digits (SAD). MFCC features ten Arabic digits spoken by $88$ speakers.
Baselines. We compare the results of several baseline methods reported by [42]. The methods cover the following paradigms: One-class classification: One-class SVM (OC-SVM), and its deep versions, DeepSVDD (“DSVDD”) [50], DROCC [24]. Tree-based detectors: Isolation Forest (IF) [37]. Density estimation: LOF, a specialized version of nearest neighbor anomaly detection [10]. DAGMM (“DAG”) [62]: density estimation in an auto-encoder latent space Auto-regressive methods - RNN and LSTM-ED (“ED”) - deep neural network-based version of auto-regressive prediction models [38]. Transformation prediction - GOAD [2] and NeuTraL-AD [42] are based on transformation prediction, and are adaptations of RotNet-based approaches (such as GEOM [22]).
Metric. Following [42], we use the series-level ROC-AUC metric.
Results. Our results are presented in Tab. 3. We can observe that different baseline approaches are effective for different datasets. $k$NN-based LOF is highly effective for SAD which is a large dataset but achieves worse results for EPSY. Auto-regressive approaches achieve strong results on CT. Transformation-prediction approaches, GOAD and NeuTraL achieve the best performance of all the baselines. The learned transformations of NeuTraL achieved better results than the random transformations of GOAD.
Our method achieves the best overall results both on average and individually on all datasets apart from SAD (where it is comparable but a little lower than NeuTraL). Note that unlike NeuTraL, our method is far simpler, does not use deep neural networks and is very fast to train and evaluate. It also has fewer hyperparameters.
@@ 5.3 Implementation Details
We provide here the main implementation details for our image anomaly detection application. Further implementation details for the image application can be found in the supplementary material (App.13.1). Implementation details for the time series experiments can be found in the supplementary material as well (App.13.2).
Multiple crops for image anomaly detection. Describing the entire image as a single set might sometimes lose discriminative power when the anomalies are localized. To mitigate this issue, we can treat only a part of an image as our entire set. To do so, we crop the image to a factor of $c$, and compare the elements taken only from these crops. We compute an anomaly score for each crop factor and for each center location. We then average over the anomaly scores of the different crop center locations for the same crop factor $c$. Finally, for each ResNet level (described above), we average the anomaly scores over the different crop ratios $c$. We use crop ratios of $\{1.0,0.7,0.5,0.33\}$. The different center locations are taken with a stride of $0.25$ of the entire image. We note that feeding an image to the network multiple provide only a marginal advantage (Tab. 5), and that baseline methods used a significantly larger of forward passes with the feature extractors [48].
@@ 5.4 Ablations
We present ablations for the image logical AD methods. For further ablations of the histogram parameters and for the time series modality, see Sec. 16.
Using individual ResNet levels. In Tab.5 we report the results when different components of our multi-level ResNet ensemble are removed. We report the results using only the representation from the third or fourth ResNet block (“Only 3 / 4”). We report the results of using both ResNet blocks but without the raw-pixels level (“No Pixels”).
@@
Ablating our histogram density-estimation method. In Tab.4 we ablate different aspects of our histogram set descriptors. Simple averaging. We show a simple averaging [35] of the set features (Fig. 2), ablating our entire set-features approach. This yields a significantly worse performance. No random projection. We ablate our use of random projections (Sec.3.3). We replace the random histograms with similar histograms using the raw given features. No whitening. We ablate our Gaussian modelling of the set features. The whitening is not essential for the image modality, as it is for the time-series data (Table 13).
@@
[…]
Only 3 Only 4 No pixels Only full Ours
Breakfa. 95.9 95.7 96.8 97.2 97.7
Juice bo. 93.0 97.0 95.8 97.0 97.1
Pushpins 79.2 67.0 74.0 89.9 88.9
Screw b. 79.8 70.4 76.6 76.2 81.1
Splicing. 84.7 85.6 86.1 90.7 91.5
Average 86.5 83.1 85.9 90.2 91.2
Density estimation with histogram ablation. We compare our method for density estimation of the elements collection as explained in Sec.5.4. We evaluate these methods using the 3rd and 4th ResNet blocks, as the raw pixels level adds significant variance over shading the difference between some of the alternatives. While ablation may give stronger results in specific cases, our set approach (instead of the feature average as in Fig.2) together with the random projections and whitening generally outperforms.
Ablating the number of bins and the number of projections. While generally we would like to have as many random projections as possible; and a large number of bins per histogram (as long we have enough statistics to estimate the occupancy in each of them) we find that in practice the values we choose are large enough. We show in the supplementary material (App. Tab.11,12) that while significantly lower values in these parameters degrade our performance, the benefit from using larger values saturates.
@@ 6 Discussion
Complementary strength of density estimation and reconstruction based approaches for logical anomaly detection. Our method and GCAD [5], a reconstruction based approach, exhibit complementary strengths. Our method is most suited to detect anomalies resulting from the distribution of featured objects in each image. E.g., object replacements, additional or missing objects, or colors indicating a logical inconsistency with the rest of the image. The generative modeling by GCAD gives stronger results when the positions of the objects are anomalous (e.g., one object containing another when it should not, or vice versa, as in the Pushpins class). The intuition here is that our approach treats the patches as an unordered set, and might not capture exact spatial relations between the objects. Therefore, it may be a natural direction to try and use both approaches together. A practical way to take advantage of both approaches would be an ensemble. Ultimately, future research is likely to lead to the development of better approaches, combining the strengths of both methods.
Relation to previous random projection methods. Our method is related to several previous methods. HBOS [23] and LODA [41] also used similar projection features for anomaly detection. Yet, these methods perform histogram-based density estimation by ignoring the dependency across projections. As they can only be applied to a single element, they do not achieve competitive performance for time series AD. Rocket/mini-rocket [18, 19] also average projection features across windows but do not tackle anomaly detection nor do they apply to image data.
Is our set descriptor approach beneficial for detecting structural image anomalies? While our method lags behind the top detection-by-segmentation approaches on structural anomalies, it achieves the top performance on specific classes. Yet, generally, detection-by-segmentation methods are better when anomalies are contained in a single element. We hypothesize that our advantage in some classes may be due to the high variation among the normal samples in these classes. In this case, too, future research may allow the construction of better detectors, enjoying the combined strength of many approaches.
Incorporating deep features for time series data. Our method can outperform the state-of-the-art in time series anomaly detection without using deep neural networks. While this is an interesting and surprising result, we believe that deep features will be incorporated into similar approaches in the future. One direction for doing this is replacing the window projection features with a suitable deep representation, while keeping the set descriptors and Gaussian modeling steps unchanged.
Fine-tuning deep features for anomaly detection. Following recent works in anomaly detection and anomaly segmentation, we used fixed pretrained features as the backbone of our method. Although some methods fine-tune deep features for anomaly detection based on the normal-only training set, we keep them constant. Doing so allows an interpretable examination of the relative strength of our novel scoring function with respect to prior works that use fixed features. Yet, we expect that fine-tuning such features could lead to further gains in the future.
@@ 7 Limitations
Element-level anomaly detection. Our method focuses on sample-level time series and image-level anomaly detection. In some applications, a user may also want a segmentation map of the most anomalous elements of each sample. We note that for logical anomalies, this is often not well defined. E.g., when we have an image with $3$ nuts as opposed to the normal $2$, each of them may be considered anomalous. To provide element-level information, our method can be combined with current segmentation approaches by incorporating the knowledge of a global anomaly (e.g., removing false positive segmentations if an image is normal). Directly applying our set features for anomaly segmentation is left for future research.

### v4

@@ 4.1 Images as Sets
Images can be seen as consisting of a set of elements at different levels of granularity. This ranges from pixels to small patches, to low-level elements such as lines or corners, up to high-level elements such as objects. For anomaly detection, we typically do not know in advance the correct level of granularity for separating between normal and anomalous samples (Heckler et al., 2023). The correct level may depend on the anomalies we will encounter, which are unknown during training. Instead, we use multiple levels of granularity, describing image patches of different sizes, and combine their scores.
In practice, we use representations from intermediate blocks of a pre-trained ResNet (He et al., 2016). As a ResNet network simultaneously embeds many local patches of each image, we pass the image samples through the network encoder and extract our representations from the intermediate activations at the end of different ResNet blocks (see Fig.3). We define each spatial location in the activation map as an element. Note that as different blocks have different resolutions, they yield different numbers of elements per layer. We run our set methods with the elements at the end of each residual block used and combine the results in an ensemble as detailed in the appendix (App.C.1).
@@ 4.2 Time Series as Sets
Time series data can be viewed as a set of temporal windows. Similarly to images, it is generally not known in advance which temporal scale is relevant for detecting anomalies. I.e., what is the duration of windows which includes the semantic phenomenon. Inspired by Rocket Dempster et al. (2020), we define the basic elements of a time series as a collection of temporal window pyramids. Each pyramid contains $L$ windows. All the windows in a pyramid are centered at the same time step, each containing $\tau$ samples (Fig.3). The first level window includes $\tau$ elements with stride $1$, the second level window includes $\tau$ elements with stride $2$, etc. Such a window pyramid is computed for each time step in the series. The entire series is represented as a set of pyramids of its elements. Implementation details for both modalities are described in Sec.C.2.
@@ 5.1 Logical Anomaly Detection Results
Logical Anomalies Dataset. We use the recently published MVTec-LOCO dataset Bergmann et al. (2022) to evaluate our method’s ability to detect anomalies caused by unusual configurations of normal elements. This dataset features five different classes: breakfast box, juice bottle, pushpins, screw bag and splicing connector (see Fig.1). Each class includes: (i) a training set of normal samples ($\sim 350$ samples). (ii) a validation set, containing a smaller set of normal samples ($\sim 60$ samples). (iii) a test set, containing normal samples, structural anomalies, and logical anomalies ($\sim 100$ each).
The anomalies in each class are divided into structural anomalies and logical anomalies. Structural anomalies feature local defects, somewhat similar to previous datasets such as Bergmann et al. (2019). Conversely, logical anomalies may violate ‘logical’ conditions expected from the normal data. As one example, an anomaly may include a different number of objects than the numbers expected from a normal sample (while all the featured object types exist in the normal class; see Fig.1). Other types of logical anomalies in the dataset may include cases where distant parts of an image must correlate with one another. For instance, within the normal data, the color of one object may correlate with the length of another object. These correlations may break in an anomalous sample.
Baselines. We compare to baseline methods used by the paper which presented the MVTec-LOCO dataset Bergmann et al. (2022): Variational Model (VM) Steger (2001), MNAD, f-AnoGAN Schlegl et al. (2017), AE / VAE. Student Teacher (ST), SPADE, PatchCore (PCore) Roth et al. (2022). We also compare to GCAD Bergmann et al. (2022) - a reconstruction-based method, based on both local and global deep ResNet features, which was explicitly designed for logical anomaly detection; EfficientAD - Batzner et al. (2023), a reconstruction-based method with a loss aimed at preventing an autoencoder from reconstructing well anomalous unseen images. We also report the results by PUAD Sugawara & Imamura (2024), an ensemble method combining Batzner et al. (2023) and Rippel et al. (2021). Finally, we report SINBAD+EfficientAD, a simple average of EfficientAD’s per-sample results and ours. As the last set of baselines does not always report per-class accuracies, we report them in a different table.
@@
Results. We report per-class results on image-level detection of logical anomalies and structural anomalies in Tab.1. Interestingly, we find complementary strengths between our approach and GCAD, a reconstruction-based approach by Bergmann et al. (2022). Although GCAD performed better on specific classes (e.g., pushpins), our approach provides better results on average. Notably, our approach provides non-trivial anomaly detection capabilities on the screw bag class, while baseline approaches are close to the random baseline performance (which is $50\%$ ROC-AUC). EfficientAD Batzner et al. (2023), focuses on structural anomalies and achieves impressive results on them, but underperforms on logical anomalies (see Tab.2).
As our method is comparatively strong on specific classes (e.g., Screw bag, Logical, $81.1\%$ compared to $56.0\%$ of GCAD and $55.5\%$ of EfficientAD), it serves as a valuable component in ensemble methods. For example, a simple combination of our method with EfficientAD Batzner et al. (2023) (SINBAD+EfficientAD) outperforms all other methods and ensembles (See table 2).
Our approach also improves upon detection-by-segmentation methods in detecting structural anomalies in some classes. This is somewhat surprising, as one may assume that detection-by-segmentation approaches would perform well in these cases. One possible reason for that is the high variability of the normal data in some of the classes (e.g., breakfast box, screw bag, Fig.1). This high variability may induce false positive detections for baseline approaches. While different methods provide complementary strengths, on average, our method provides state-of-the-art results in logical anomaly detection. See also the discussion at Sec.6.
@@
See Tab.7 for the complete table with error bounds.
@@
5.2 Time Series anomaly Detection Results
Time series dataset. We evaluate on the five datasets used in NeurTraL-AD Qiu et al. (2021): RacketSports (RS). Accelerometer and gyroscope recordings of players playing different racket sports. Each sport is designated as a class. Epilepsy (EPSY). Accelerometer recording of healthy actors simulating four activity classes, e.g. an epileptic shock. Naval air training and operating procedures standardization (NAT). Positions of sensors mounted on body parts of a person performing activities. There are six different activity classes in the dataset. Character trajectories (CT). Velocity trajectories of a pen on a tablet. There are $20$ characters in this dataset. Spoken Arabic Digits (SAD). MFCC features ten Arabic digits spoken by $88$ speakers.
Baselines. We compare the results of several baseline methods reported by Qiu et al. (2021). The methods cover the following paradigms: One-class classification: One-class SVM (OC-SVM), and its deep versions, DeepSVDD (“DSVDD”) Ruff et al. (2018), DROCC Goyal et al. (2020). Tree-based detectors: Isolation Forest (IF) Liu et al. (2008). Density estimation: LOF, a version of nearest neighbor anomaly detection Breunig et al. (2000). DAGMM (“DAG”) Zong et al. (2018): density estimation in an auto-encoder latent space. Auto-regressive methods - RNN and LSTM-ED (“ED”) - deep neural network-based version of auto-regressive prediction models Malhotra et al. (2016). Transformation prediction - GOAD Bergman & Hoshen (2020) and NeuTraL-AD Qiu et al. (2021) are based on transformation prediction, and are adaptations of RotNet-based approaches such as GEOM Golan & El-Yaniv (2018).
Metric. Following Qiu et al. (2021), we use the series-level ROC-AUC metric.
Results. Our results are presented in Tab. 3. We can observe that different baseline approaches are effective for different datasets. $k$NN-based LOF is highly effective for SAD which is a large dataset but achieves worse results for EPSY. Auto-regressive approaches achieve strong results on CT. Transformation-prediction approaches, GOAD and NeuTraL achieve the best performance of all the baselines. The learned transformations of NeuTraL achieved better results than the random transformations of GOAD. Our method achieves the best overall results both on average and individually on all datasets apart from SAD, where it is comparable but a little lower than NeuTraL. We note that unlike NeuTraL, our method is far simpler, does not use deep neural networks, and is very fast to train and evaluate. It also has fewer hyperparameters.
@@ 5.3 Implementation Details
We provide here the main implementation details for our image anomaly detection application. Further implementation details for the image application can be found in App.C.1; Implementation details for the time series application can be found in the supplementary material in App.C.2.
ResNet levels. We use the representations from the 3rd and 4th blocks of a WideResNet50$\times$2 (resulting in sets size $7\times 7$ and $14\times 14$ elements, respectively). We also use all the raw pixels in the image as an additional set (resized to $224\times 224$ elements). The total anomaly score is the average of the anomaly scores obtained for the set of 3rd ResNet block features, the set of 4th ResNet block features, and the set of raw pixels. The average anomaly score is weighted by the following factors $(1,1,0.1)$ respectively (see App.D for our robustness to the choice of weighting factor).
Multiple crops for image anomaly detection. Describing the entire image as a single set might sometimes lose discriminative power when the anomalies are localized. To mitigate this issue, we can treat only a part of an image as our entire set. To do so, we crop the image to smaller images by a factor of $c$, and compare the elements taken from each crop. We compute an anomaly score for each crop scale and for each center location. We then average over the anomaly scores of the different crop center locations for the same crop scale $c$. Finally, for each ResNet level, we average the anomaly scores over the different crop scales. We use crop scales of $\{1.0,0.7,0.5,0.33\}$. The different center locations are taken with a stride of $0.25$ of the entire image. We observe that combining different scales offers only a marginal advantage (see Tab.4), while baseline methods require significantly more forward passes through the feature extractors Roth et al. (2022).
@@ 5.4 Ablations
We present ablations for the image logical AD methods. For further ablations of the histogram parameters and for the time series application, see appendix F.
Using individual ResNet levels. In Tab.4 we report the results when different components of our multi-level ResNet ensemble are removed. We report the results using only the representation from the third or fourth ResNet block (“Only 3” / “Only 4”). We also report the results using both ResNet blocks but without the raw-pixels level (“No Pixels”).
@@
Ablating our histogram density-estimation method. In Tab.5 we ablate different aspects of our histogram set descriptors. Sim. Avg. We show a simple averaging Lee et al. (2018) of the set features (see also Fig. 2), ablating our entire set-features approach. This yields a significantly worse performance. No proj. We ablate our use of random projections (Sec.3.3). We replace the random histograms with similar histograms using the raw given features. No whit. We ablate our Gaussian modeling of the set features. The whitening is not essential for the image modality, as it is for the time-series data (see Tab.12). We compare these variants of our method using the 3rd and 4th ResNet blocks, as the raw pixels level adds significant variance overshadowing the difference between some of the alternatives. While ablation may give stronger results in specific cases, our set approach together with the random projections and whitening generally outperforms.
[…]
Only 3 Only 4 No pixels Only full Ours
Breakfast box 95.9 95.7 96.8 97.2 97.7
Juice bottle 93.0 97.0 95.8 97.0 97.1
Pushpins 79.2 67.0 74.0 89.9 88.9
Screw bag 79.8 70.4 76.6 76.2 81.1
Splicing connectors 84.7 85.6 86.1 90.7 91.5
Average 86.5 83.1 85.9 90.2 91.2
@@
Ablating the number of bins and the number of projections. While generally we would like to have as many random projections as possible; and a large number of bins per histogram, as long we have enough statistics to estimate the occupancy in each of them; We find that in practice the values we choose are large enough. We show in the App. Tab.10,11 that while significantly lower values in these parameters degrade our performance, the benefit from using larger values saturates.
@@ 6 Discussion
Complementary strength of density estimation and reconstruction based approaches for logical anomaly detection. Our method and GCAD Bergmann et al. (2022), a reconstruction-based approach, exhibit complementary strengths. Our method is most suited to detect anomalies resulting from the distribution of featured objects in each image. E.g., object replacements, additional or missing objects, or components indicating a logical inconsistency with the rest of the image. The generative modeling by GCAD gives stronger results when the positions of the objects are anomalous. E.g., one object containing another when it should not, or vice versa, as in the Pushpins class. The intuition here is that our approach treats the patches as an unordered set, and might not capture exact spatial relations between the objects. Therefore, it may be a natural direction to try and use both approaches together. A practical way to take advantage of both approaches would be an ensemble as we suggest in Sec.5. Ultimately, future research is likely to lead to the development of better approaches, combining the strengths of both methods.
Relation to previous random projection methods. Our method is related to several previous methods. HBOS Goldstein & Dengel (2012) and LODA Pevnỳ (2016) also used similar projection features for anomaly detection. Yet, these methods perform histogram-based density, ignoring the dependency across projections. As they can only be applied to a single element, they do not achieve competitive performance for time series AD. Rocket and mini-rocket Dempster et al. (2020; 2021) also average projection features across selected windows from a given sample but do not apply to image data.
Additional discussion points can be found in App.H.
@@ 7 Limitations
Detecting structural anomalies. Our approach aims to detect specific, yet important, types of anomalies - image-level logical anomalies and the analogues time-series sequence-level anomalies. It is not particularly effective for detecting local structural anomalies, such as scratches or dents in images of objects Bergmann et al. (2019); Zou et al. (2022), object-level anomalies Reiss et al. (2021), or local time-series anomalies Blázquez-García et al. (2021). We evaluate our method on appropriate datasets. Currently, only one dataset evaluates logical anomalies Bergmann et al. (2022). Yet, this very comprehensive dataset contains 5 different sub-tasks, where each sub-task features numerous different types of anomalies. An anomaly detection method must rely on some assumptions regarding the nature of the anomalies one wishes to discover (Reiss et al., 2023). Therefore, when the type of anomalies is unknown, we recommend combining our method with methods tailored to different types of anomalies (Reiss et al., 2022; Roth et al., 2022; Batzner et al., 2023).
Element-level anomaly detection. Our method focuses on sequence-level time series and image-level anomaly detection. In some applications, a user may also want a segmentation map of the most anomalous elements of each sample. We note that for logical anomalies, this is often not well defined. E.g., when we have an image with $3$ nuts as opposed to the normal $2$, each of them may be considered anomalous. To provide element-level information, our method can be combined with current segmentation approaches by incorporating the knowledge of a global anomaly. E.g., removing false positive segmentations if an image is normal. Directly applying our set features for anomaly segmentation is left for future research.